Mole Fraction 1. Calculate the mole fraction of each component of a solution containing 65 g of ethanol (C₂H6O) in 350g of water.

1
Expert's answer
2022-03-12T02:00:24-0500

Solution:

solute = ethanol (C2H6O)

solvent = water (H2O)


The molar mass of ethanol (C2H6O) is 46.07 g/mol

The molar mass of water (H2O) is 18.015 g/mol


Calculate the moles of each component:

Moles of C2H6O = (65 g C2H6O) × (1 mol C2H6O / 46.07 g C2H6O) = 1.411 mol C2H6O

Moles of H2O = (350 g H2O) × (1 mol H2O / 18.015 g H2O) = 19.428 mol H2O

Therefore,

Total moles of solution = Moles of C2H6O + Moles of H2O = 1.411 mol + 19.428 mol = 20.839 mol


Calculate the mole fraction (χ) of each component:

χ = Moles of component / Total moles of solution

Therefore,

χsolute = χC2H6O = Moles of C2H6O / Total moles of solution = (1.411 mol) / (20.839 mol) = 0.0677

χC2H6O = 0.0677

χsolvent = χH2O = Moles of H2O / Total moles of solution = (19.428 mol) / (20.839 mol) = 0.9323

χH2O = 0.9323


Answer:

The mole fraction of ethanol is 0.0677

The mole fraction of water is 0.9323